OpenStudy (honeymonster):

can you check this answer please..i'm not sure if its correct :S i'll post it in replies

OpenStudy (honeymonster):

\[3(y+2) = 2(y-1)\] 3y+6=2y-2 3y-2y=-2-6 y=-8

OpenStudy (anonymous):

looks good to me

hartnn (hartnn):

correct :)

OpenStudy (cathyangs):

That's perfect! :)

OpenStudy (honeymonster):

awesome..thank you :)